Determine the shift in the wavelength (in m) of scattered x-rays that are observed at an angle of 44.3° to the incident beam (in m).
Answer #1

given

theta = 44.3 degrees

we know,
shift in the wavelength,

lamda' - lamda = (h/(m*c))*(1 - cos(theta)


= (6.625*10^-34/(9.1*10^-31*3*10^8))*(1 - cos(44.3))

= 6.90*10^-13 m <<<<<<------------Answer